# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ 2.4 Primary and secondary catalysis
Based on this analysis, we found that disruptions of primary edges as a result of a knockout were significantly associated with greater fitness losses than disruptions of secondary edges (Wilcoxon rank-sum test, $p<10^{-6}$, Figure 3D). In addition, primary disruptions that were only partial were associated with more contained fitness losses. Finally, minor fitness gains could be identified when mutants were associated with secondary and partial primary disruptions, but not when complete primary disruptions occurred. We did not find as significant differences between complete disruptions and full primary disruptions (Supplementary Figure 8), supporting the idea that catalytic relationships annotated as secondary are unlikely to be strong enough to stand-in for disrupted primary ones under normal physiological conditions.
@@ 2.5 Enzyme-constrained Flux Balance Analysis with EC-iCH360
To extend the scope and applicability of iCH360, we constructed a version of the model that allows for enzyme-constrained flux simulations, which we denote as EC-iCH360. We constructed EC-iCH360 in the sMOMENT [bekiaris_automatic_2020] format (Methods). The sMOMENT framework is inherently simple and generates the same solution space as more complex model formats, such as GECKO, unless reaction-specific capacity constraints are specified in the latter [sanchez_improving_2017, bekiaris_automatic_2020]. Since it requires unique reaction-enzyme mappings, we used the knowledge graph to remove all secondary catalytic relationships in the model (Section 2.4).
@@
We first parametrized the model by defining a flux-specific enzyme cost for each reaction (in units of grams of enzyme per unit flux), using as values the estimated in vivo turnover number from [heckmann_kinetic_2020]. Using this enzyme-constrained model, we then predicted enzyme abundances and compared them to experimental enzyme abundances, estimated from a dataset of proteomic measurements for aerobic growth on eight different carbon sources [schmidt_quantitative_2016](Methods). This analysis led to generally good predictions, with root mean squared error (RMSE, computed for log10-transformed enzyme abundances) ranging from $0.53$ to $0.62$ (Supplementary Figure 9). In order to assess the nature of residuals between measurements and predictions, we investigated the geometric mean of enzyme abundances across all eight conditions (Supplementary Figure 9, bottom right panel). Notably, averaging predicted and measured enzyme abundances, respectively, across conditions did not significantly improve the RMSE, indicating that the abundances of individual enzymes were systematically over- or under-predicted across conditions.
@@
To increase the predictive ability of the model, we thus adjusted the turnover numbers by fitting them to experimental measurements through a custom heuristic (Methods, Supplementary Information A.5). By simultaneously fitting all available conditions, we ensured that our adjustment procedure is robust to condition-specific biases. Further, by introducing regularisation within our adjustment scheme, we penalised large deviations of parameters from the original dataset, increasing the robustness of the procedure to overfitting. Our adjusted parameter set shows a mean absolute deviation (computed for log10-transformed turnover values) from the original parameter set of $\approx 0.22$ (Supplementary Figure 10) and results in significantly better agreement with experimental measurements across conditions (Figure 4A). Further, mean enzyme abundances across conditions are very well predicted with the adjusted parameter set (Figure 4B), implying that residuals between measurements and predictions are now to be attributed to variability in enzyme demand across conditions (which simple frameworks such as enzyme-constrained FBA cannot account for), rather than systematic over- or under-predictions. Thus, each adjusted turnover parameter can be thought of as a ”typical” apparent $k_{\rm cat}$ value, incorporating average saturation trends for an enzyme across growth conditions.
@@ 2.6 Elementary Flux Modes in the reduced model variant iCH360red
Despite the small size of iCH360, we found the explicit enumeration of its elementary flux modes (EFMs) to be intractable. This is not necessarily surprising, since the topology of a metabolic network is a much better determinant of EFM count than its sheer size. Metabolic networks can possess different types of redundancies, such as the presence of alternative pathways for the production of the same metabolite, the use of alternative cofactors for the same catalytic step in a pathway, or the presence of alternative transporters for the uptake/excretion of a compound. While knowledge about these redundancies is often valuable, including them in the model can increase the number of EFMs exponentially, hampering or even preventing EFM-based analyses.
@@
To address this issue, we identified and removed a small set of such redundancies in iCH360, using available information from literature, whenever possible, to ensure the most physiologically relevant alternative was maintained (Supplementary Table 3). This resulted a metabolic submodel of iCH360, a model variant that we denote by iCH360red. iCH360red contains $305$ metabolic reactions ($18$ less than iCH360) and shares the same production envelopes with its parent model for a number of metabolites of interest (Supplementary Figure 12). While the number of EFMs in iCH360red is still relatively large ($\approx 13.5$ millions for aerobic growth on glucose, see Supplementary Table 4), it is not prohibitive for most types of EFM-based analysis, and their explicit enumeration does not require high-performance computing (Methods).
@@
We used the obtained EFMs to study the possible combinations of achievable growth rates and yields in the network [wortel_metabolic_2018]. To this end, we considered growth on glucose as a scenario and computed, for each EFM from iCH360red, its yield and its achievable cell growth rate, which we estimate based on the enzyme costs defined for the enzyme-constrained model (Methods). Based on this analysis, we identified a front of Pareto-optimal EFMs, along which any increase in the growth rate will necessarily lead to a reduction in yield (Figure 5). Along the Pareto front, we observe a transition from a purely respiratory mode at maximum yield (Supplementary Figure 13) to a mixed respiratory-fermentative mode at maximum growth (Supplementary Figure 14). Quantitatively, the extent of this tradeoff was rather modest: the EFM with maximal yield reaches almost the maximal growth rate, so only minor gains in growth rate can be achieved by using other, fermentative modes along the Pareto front. However, it is worth noting that this analysis was performed using a simple capacity-based enzyme cost function, which summarises different determinants of enzyme cost such as driving force and substrate saturation [noor_protein_2016] in one single value, assumed to be constant. Hence, repeating the analysis with a more complete enzyme-cost function [noor_protein_2016] could help elucidate the nature of this tradeoff. To demonstrate that the shape and size of the Pareto front depend strongly on growth conditions, we also simulated an environment with very low oxygen levels. We implemented this by increasing the flux-specific enzyme cost of the oxygen-dependent reaction in the respiratory chain, equivalent to assuming a lower enzyme efficiency due to a lower oxygen level (Methods). Results (Supplementary Figure 15) show a much broader front of Pareto optimal EFMs, indicating that the nature of the observed tradeoff is indeed condition-dependent. Notably, a similar dependence of the Pareto front on extracellular oxygen availability has been previously observed in a small-scale model of E. coli core metabolism [wortel_metabolic_2018].
@@ 2.7 Saturation FBA and modelling of overflow metabolism
In order to study the effect of external conditions on optimal metabolic strategies in more detail, we used another framework that does not require an enumeration of EFMs and that allows for additional flux bounds, for example, a lower bound on the ATP consumption for cell maintenance. The saturation FBA (satFBA) framework [muller_resource_2015] is a variant of enzyme-constrained FBA, wherein a fixed enzyme cost per flux is assumed for all metabolic reactions in a model, except for the substrate transporter, for which a complete kinetic law is used (Methods). Since the external substrate concentration is a simple parameter, screening this concentration is equivalent to screening the values of the transporter efficiency. Here we used satFBA to simulate how the growth-maximising solution of the network varies in response to changing extracellular glucose concentration. By solving the satFBA problem for a range of glucose concentrations, we predicted the dependence of the cell’s growth rate on substrate concentration, resulting in the typically observed Monod curve (Figure 6A). If the problem does not contain any further flux bounds (so that the magnitude of fluxes at the optimum is solely limited by the maximum enzyme availability) the solution of satFBA problems will be an elementary flux mode [muller_resource_2015]. Hence, in this case we can use satFBA to explore how a cell should switch across elementary modes as a function of the growth environment.
@@
At low glucose concentrations, the glucose transporter operates at low saturation and glucose uptake is enzymatically expensive, leading to a high-yield, purely respiratory metabolic mode at the optimum (Figure 6B and C). As substrate availability is increased, the cost of substrate uptake decreases and higher growth rates are achieved by switching to lower yield, acetate-secreting modes, in line with typical observations of overflow metabolism in E. coli [basan_overflow_2015, chen_energy_2019]. However, the satFBA formalism can also be used with additional flux bounds, thus going beyond EFM-based analysis. For example, if a positive lower bound on ATP hydrolysis is added as a maintenance requirement, optimal solutions to the satFBA problem will no longer be elementary modes, and the yield of the optimal solution no longer follows a piecewise constant profile (Supplementary Figure 16)
@@
Figure 6: Saturation FBA enables the exploration of the optimal switching across elementary modes as a function of the growth environment. A: satFBA predictions for the growth rate as a function of external glucose concentration, showing a typical Monod curve. Note that, satFBA computes the cell’s growth rate by assuming a fixed total enzyme mass budget while varying the saturation of the substrate transporter as a function of external substrate concentration. Importantly, although the curve is continuous and smooth, it comprises many smaller sections where each one has a different dominating elementary mode. B: satFBA predictions for the acetate excretion flux, showing progressively higher use of fermentative metabolism in the optimal solution as external glucose availability increases. C: The yield of the optimal satFBA solution, computed as the ratio of biomass flux to glucose uptake, progressively decreases as external glucose availability increases.
@@ 2.8 Equilibrium constants, thermodynamic forces, and thermodynamically feasible states
Living systems operate out of thermodynamic equilibrium, and thermodynamics places strong constraints on the operation of metabolic systems. In any metabolic state, the flux directions must follow the signs of thermodynamic forces, which depend on metabolite concentrations and equilibrium constants. To provide these constants as parts of our model, we used the component contribution framework [noor_consistent_2013] to estimate the standard Gibbs free energy ($\Delta_{r}G^{\prime\circ}$) of each reaction (Methods). These estimates account for compartment-specific chemical environment parameters, such as pH, pMg, and ionic strength, and were corrected to account for protons and charge translocation in multi-compartment reactions.
@@
The resulting parameter set covers the vast majority of the model (over 97 % of metabolic reactions, with the remaining being not covered by the component contribution database used here) and accounts for the uncertainty in the estimates through a multivariate covariance matrix. Accounting for correlations between the different $\Delta_{r}G^{\prime\circ}$ values becomes important when imposing thermodynamic constraints on the model. For example, the fatty acid biosynthesis subsystem in the model consists of repeated elongation cycles, where a short sequence of chemical transformations is repeatedly performed on a growing carbon chain. As a result, even if the $\Delta_{r}G^{\prime\circ}$ for each reaction in the pathway is known with some uncertainty, this uncertainty is tightly correlated across the reactions, which constrains the set of achievable thermodynamic states in the network.
@@
Using this set of thermodynamic constants, we first tested whether some typical flux distributions obtained from the model are thermodynamically feasible. To this end, we considered flux distributions generated by parsimonious Flux Balance Analysis (pFBA) across $12$ growth conditions and computed their max-min driving force (MDF) [noor_pathway_2014], accounting for uncertainty in the estimates (Methods). We found a positive MDF for each of the flux distributions, indicating that all pFBA solutions tested are thermodynamically feasible. Notably, we found that the computed MDF values cluster very clearly in three groups, corresponding to aerobic growth on glycolytic substrates, aerobic growth on gluconeogenic substrates, and anaerobic conditions (Figure 7A).
@@
Having confirmed that our reference FBA-derived flux distributions are thermodynamically feasible, we then employed an alternative flux prediction method that ensures thermodynamically feasibilty by construction. The probabilistic metabolic optimisation (PMO) framework [gollub_probabilistic_2021] uses a mixed-integer quadratic programming approach to compute a set of fluxes, metabolite concentrations, and reaction driving forces which is probabilistically most in agreement with experimentally measured metabolite concentrations (Methods). Using this framework, we computed a maximum-likelihood thermodynamic state for the model, simulating aerobic growth on glucose. In the thermodynamic state computed by PMO, we found that all metabolites lie within physiologically reasonable concentrations (1 $\mu$M – 1 mM). In addition, all ”anomalous concentrations” identified by the framework (metabolite concentrations lying more than one standard deviation away from the mean experimental value) had been identified and explained previously [gollub_probabilistic_2021], and can potentially be addressed by lumping together those reactions for which substrate channelling is known to happen in literature.
@@
We used the PMO-derived thermodynamic state to identify candidate bottlenecks, in terms of enzyme demand, across the network. To this end, we first computed the flux-force efficacy of each reaction in the thermodynamic state [noor_pathway_2014] (Methods). The flux-force efficacy is a unitless quantity, ranging from 00 to $1$, denoting the ratio between net flux (forward minus backward flux) and total flux (forward plus backward flux) of a reaction. Reactions operating at low flux-force efficacy have a lower net flux due to two reasons: (1) the forward flux is lower in absolute terms, and (2) the higher backward flux is counter-productive and subtracts from the forward flux. Therefore, to achieve a given required net flux, the cell has to invest more resources in maintaining a higher enzyme level. To identify potential thermodynamic bottlenecks leading to high enzyme costs in specific reactions, we thus screened all reactions for their predicted flux-force efficacy and flux (Figure 7B) and identified those predicted to operate at low efficacy, while carrying significant flux (Figure 7B, labelled points). To assess the extent to which the thermodynamic operating states predicted by PTA are predictive of enzyme investment, we split reactions in two groups based on whether their predicted flux-force efficacy sits above or below $50\%$, and compared the distributions of measured enzyme abundances between the two groups (Methods, Supplementary Figure 17). While there are many determinants of enzyme abundance beyond thermodynamics, including enzyme turnover, affinity, or regulation, we observed a significant difference between the two distributions ($p<0.001$, two-sided Wilcox sum-rank test), with the low efficacy having approximately a 3-fold higher median enzyme abundance than the high efficacy group.
@@
Figure 7: Thermodynamic analysis of the model via the curated thermodynamic parameter set. A Probabilistic Min-max driving force analysis of flux distributions obtained by parsimonious flux balance analysis for a total of $12$ growth conditions. All flux distributions tested have a positive MDF, implying they are thermodynamically feasible under reasonable physiological metabolite concentration ranges. The computed MDF values cluster in three groups, corresponding to glycolytic aerobic, gluconeogenic aerobic, and anaerobic growth conditions. B: Fluxes relative to the glucose uptake flux (EX_glc__D_e) and flux-force efficacies computed by probabilistic metabolic optimisation (PMO). The labelled data points represent examples of reactions (excluding transport and spontaneous reactions) with low predicted flux-force efficacy (here, below $20\%$), but carrying high relative flux in the optimal solution (here, more than $5\%$ of the glucose uptake flux). xyl__D: D-xylose; glc__D: D-glucose; rib__D: D-ribose; glyc: glycerol; akg: alpha-keto-glutarate; ac: acetate; pyr: pyruvate; lac__D: D-lactate; succ: succinate. fum: fumarate

### v3

@@ 2.3 Connecting reactions to their catalysing enzymes and the enzymes’ protein components
Based on this analysis, we found that disruptions of the primary edges as a result of a knockout were significantly associated with greater fitness losses than disruptions of secondary edges (Wilcoxon rank-sum test, $p<10^{-6}$, Figure 3D). In addition, primary disruptions that were only partial were associated with more contained fitness losses. Finally, minor fitness gains could be identified when mutants were associated with secondary and partial primary disruptions, but not when complete primary disruptions occurred. We did not find any significant differences between complete disruptions and full primary disruptions (Supplementary Figure S10), supporting the idea that catalytic relationships annotated as secondary are unlikely to be strong enough to stand-in for disrupted primary ones under normal physiological conditions.
@@ 2.4 Enzyme-constrained Flux Balance Analysis with EC-iCH360
To make iCH360 applicable for enzyme-constrained flux simulations, we constructed a version of the model containing all necessary extra information, which we denote as EC-iCH360. We constructed EC-iCH360 in the sMOMENT [bekiaris_automatic_2020] format (Methods). The sMOMENT framework is inherently simple and generates the same solution space as more complex model formats, such as GECKO, unless reaction-specific capacity constraints are specified in the latter [sanchez_improving_2017, bekiaris_automatic_2020]. Since it requires unique reaction-enzyme mappings, we used the knowledge graph to remove all secondary catalytic relationships in the model (Section 2.3).
@@
We first parametrised the model by defining a flux-specific enzyme cost for each reaction (in units of grams of enzyme per unit flux), using as values the estimated in vivo turnover numbers from [heckmann_kinetic_2020]. Using this enzyme-constrained model, we then predicted enzyme abundances and compared them to experimental enzyme abundances, estimated from a data set of proteomic measurements for aerobic growth on eight different carbon sources [schmidt_quantitative_2016] (Methods). This analysis led to predictions with root mean squared error (RMSE, computed for log10-transformed enzyme abundances) ranging from $0.53$ to $0.62$ (Supplementary Figure S11). To assess the nature of residuals between measurements and predictions, we investigated the geometric mean of enzyme abundances across all eight conditions (Supplementary Figure S11, bottom right panel). If the mismatch between measurements and predictions were due to each enzyme operating at different saturation levels in each condition, one would expect that averaging would reduce these differences. However, averaging the predicted and measured enzyme abundances, respectively, across conditions did not significantly improve the RMSE, indicating that the abundances of individual enzymes were systematically over- or under-predicted between conditions.
@@
To increase the predictive capacity of the model, we adjusted the turnover numbers by fitting them to experimental measurements through a custom heuristic (Methods, Supplementary information A.5). By simultaneously fitting all available conditions with a single set of parameters, we ensured that our adjustment procedure is robust to condition-specific biases. Furthermore, by introducing regularisation within our adjustment scheme, we penalised large deviations of parameters from the original data set, increasing the robustness of the procedure to overfitting. Our adjusted parameter set shows a mean absolute deviation (computed for log10-transformed turnover values) from the original parameter set of $\approx$ 0.22 (Supplementary Figure S12) and results in significantly better agreement with experimental measurements across conditions (Figure 4A). Further, mean enzyme abundances across conditions are very well predicted with the adjusted parameter set (Figure 4B), implying that residuals between measurements and predictions are now to be attributed to variability in enzyme saturation across conditions (which simple frameworks such as enzyme-constrained FBA, which use a constant enzyme cost per unit flux, cannot account for), rather than systematic over- or under-predictions. Thus, each adjusted turnover parameter can be thought of as a “typical” apparent $k_{\rm cat}$ value, incorporating average saturation trends for an enzyme across growth conditions. To further confirm this aspect, we performed a leave-one-out cross-validation analysis (Methods), where each condition was in turn excluded from the model fitting dataset, but used to evaluate predictions. Results (Supplementary Figure S14) show that the predictions of enzyme abundances in a given condition are considerably improved even when data from that condition is not used for fitting, confirming that our parameter fitting heuristics is capturing global trends and not condition-specific effects. Training the model on the full dataset, as we did to compute the final parameter set, further improves predictions, even if to a lesser extent. This can be explained by noting that, by design, our procedure cannot adjust the turnover parameters for enzymes associated with zero flux in the reference flux distribution used for fitting (Supplementary information A.5). Hence, including all conditions in the training set allows for the turnover number of highly condition-specific enzymes (those associated to nonzero flux in only one of the reference flux distributions used by the procedure) to also be adjusted, further improving the overall prediction metrics.
@@ 2.5 Elementary Flux Modes in the reduced model variant iCH360red
Despite the small size of iCH360, we found the explicit enumeration of its elementary flux modes (EFMs) to be intractable. This is not necessarily surprising, since EFM count is crucially dependent on the topology of a metabolic network rather than its sheer size. Metabolic networks can possess different types of redundancy, such as the presence of alternative pathways for the production of the same metabolite, the use of alternative cofactors for the same catalytic step in a pathway, or the presence of alternative transporters for the uptake/excretion of a compound. Although knowledge about these redundancies is often valuable, including them in the model can increase the number of EFMs exponentially, hampering or even preventing EFM-based analyses.
@@
To address this issue, we identified and removed a small set of alternative metabolic routes in iCH360, using available information from the literature, whenever possible, to ensure that the most physiologically relevant alternative was maintained (Supplementary Table 4). This resulted in a metabolic submodel of iCH360, a model variant that we denote by iCH360red. iCH360red contains $305$ metabolic reactions ($18$ less than iCH360) and shows the same production envelopes as its parent model for a number of metabolites of interest (Supplementary Figure S15). While the number of EFMs in iCH360red is still relatively large ($\approx 13.5$ millions for aerobic growth on glucose, see Supplementary Table 5), it is not prohibitive for most types of EFM-based analysis, and their explicit enumeration does not require high-performance computing (Methods).
@@
We used the EFMs of iCH360red to study the possible combinations of achievable growth rates and yields in the network [wortel_metabolic_2018]. To this end, we considered growth on glucose as a scenario and computed, for each EFM from iCH360red, its yield, computed as the ratio of biomass flux and glucose uptake, and its achievable cell growth rate, which we estimated based on the enzyme costs defined for the enzyme-constrained model (Methods). Based on this analysis, we identified a front of Pareto-optimal EFMs, along which any increase in the growth rate will necessarily lead to a reduction in yield (Figure 5). Along the Pareto front, we observe a transition from a purely respiratory mode at maximum yield (Supplementary Figure S16) to a mixed respiratory-fermentative mode at maximum growth (Supplementary Figure S17). Quantitatively, the extent of this trade-off was rather modest: the EFM with maximal yield reaches almost the maximal growth rate, so only minor gains in growth rate can be achieved by using other, fermentative modes along the Pareto front. However, it is worth noting that this analysis was performed using a simple capacity-based enzyme cost function, which ignores metabolite concentrations by assuming a constant enzyme cost per unit flux for each reaction. Repeating the analysis with a more complete enzyme-cost function, such as one that accounts for variable thermodynamic driving force and enzyme saturation, could help elucidate the nature of this trade-off [noor_protein_2016]. To demonstrate that the shape and size of the Pareto front depend strongly on growth conditions, we also simulated an environment with very low oxygen levels. We implemented this by increasing the flux-specific enzyme cost of the oxygen-dependent reaction in the respiratory chain, equivalent to assuming a lower enzyme efficiency due to a lower oxygen level (Methods). Results (Supplementary Figure S18) show a much broader front of Pareto optimal EFMs, indicating that the nature of the observed trade-off is indeed condition-dependent. Notably, a similar dependence of the Pareto front on extracellular oxygen availability has been previously observed in a small-scale model of E. coli core metabolism [wortel_metabolic_2018].
@@ 2.6 Saturation FBA and modelling of overflow metabolism
In order to study the effect of external conditions on optimal metabolic strategies in more detail, we used another framework which does not require an enumeration of EFMs and allows for additional flux bounds, for example, to impose a minimum ATP consumption rate for cell maintenance. The saturation FBA (satFBA) framework [muller_resource_2015] is a variant of enzyme-constrained FBA, wherein a fixed enzyme cost per flux is assumed for all metabolic reactions in a model, except for the substrate transporter, for which a complete kinetic law is used (Methods). Since the external substrate concentration is a simple parameter, screening this concentration is equivalent to screening the values of the transporter efficiency. Here we used satFBA to simulate how the growth-maximising solution of the network varies in response to changing extracellular glucose concentration. By solving the satFBA problem for a range of glucose concentrations, we predicted the dependence of the cell’s growth rate on substrate concentration, resulting in the typically observed Monod curve (Figure 6A). If the problem does not contain any further flux bounds (so that the magnitude of fluxes at the optimum is solely limited by the maximum enzyme availability) the solution of satFBA problems will be an elementary flux mode [muller_resource_2015, wortel_metabolic_2014]. Hence, in this case we can use satFBA to explore how a cell should switch across elementary modes as a function of the growth environment.
@@
At low glucose concentrations, the glucose transporter operates at low saturation and glucose uptake is enzymatically expensive, leading to a high-yield, purely respiratory metabolic mode at the optimum (Figure 6B and C). As substrate availability is increased, the cost of substrate uptake decreases and higher growth rates are achieved by switching to lower yield, acetate-secreting modes [muller_resource_2015]. Since the yield of a flux distribution is, by definition, constant along an elementary flux mode, the yield varies in step-like manner as a function of external glucose concentration (Figure 6C, inset), where each jump represents a change of optimal mode. The satFBA formalism can also be used with additional flux bounds. For example, if a positive lower bound on ATP hydrolysis is added as a maintenance requirement, optimal solutions to the satFBA problem will no longer be elementary modes, and the yield of the optimal solution no longer follows a piecewise constant profile (Supplementary Figure S19C).
@@
Figure 6: Saturation FBA enables the exploration of the optimal switching across elementary modes as a function of the growth environment. A: satFBA predictions for the growth rate as a function of external glucose concentration, showing a typical Monod curve. Note that satFBA computes the cell’s growth rate by assuming a fixed total enzyme mass budget while varying the saturation of the substrate transporter as a function of external substrate concentration. Importantly, although the curve is continuous and smooth, it comprises many smaller sections, each dominated by a different elementary mode. B: satFBA predictions for the acetate excretion flux, showing progressively higher use of fermentative metabolism in the optimal solution as external glucose availability increases. C: The yield of the optimal satFBA solution progressively decreases in a step-like manner as external glucose availability increases. Each jump represents a switch in the optimal elementary flux mode.
@@ 2.7 Equilibrium constants, thermodynamic forces, and thermodynamically feasible states
Living systems operate outside of thermodynamic equilibrium, and thermodynamics places strong constraints on the operation of metabolic systems. In any metabolic state, the flux directions must follow the signs of thermodynamic forces, which depend on metabolite concentrations and equilibrium constants. To provide these constants as parts of our model, we used the component contribution framework [noor_consistent_2013] to estimate the standard Gibbs free energy ($\Delta_{r}G^{\prime\circ}$) of each reaction (Methods). These estimates account for compartment-specific chemical environment parameters, such as pH, pMg, and ionic strength, and were corrected to account for protons and charge translocation in multi-compartment reactions.
@@
The resulting parameter set covers the vast majority of the model (over 97% of metabolic reactions, with the remaining being not covered by the component contribution database used here) and accounts for the uncertainty in the estimates through a multivariate covariance matrix. Accounting for the correlations between different $\Delta_{r}G^{\prime\circ}$ values becomes important when imposing thermodynamic constraints on the model. For example, the fatty acid biosynthesis subsystem in the model consists of repeated elongation cycles, where a short sequence of chemical transformations is repeatedly performed on a growing carbon chain. As a result, even if the $\Delta_{r}G^{\prime\circ}$ for each reaction in the pathway is known with some uncertainty, this uncertainty is tightly correlated across the reactions, which constrains the set of achievable thermodynamic states in the network.
@@
Using this set of thermodynamic constants, we first tested whether some typical flux distributions obtained from the model are thermodynamically feasible. To this end, we considered flux distributions generated by parsimonious Flux Balance Analysis (pFBA) across $12$ growth conditions and computed their max-min driving force (MDF) [noor_pathway_2014], accounting for uncertainty in the estimates (Methods). We found a positive MDF for each of the flux distributions, indicating that all pFBA solutions tested are thermodynamically feasible. Notably, we found that the computed MDF values cluster very clearly in three groups, corresponding to aerobic growth on glycolytic substrates (high MDF), aerobic growth on gluconeogenic substrates (medium MDF), and anaerobic conditions (low MDF, Figure 7A).
@@
Having confirmed that our reference FBA-derived flux distributions are thermodynamically feasible, we then employed an alternative flux prediction method that ensures thermodynamic feasibility by construction. The probabilistic metabolic optimisation (PMO) framework [gollub_probabilistic_2021] uses a mixed-integer quadratic programming approach to compute a set of fluxes, metabolite concentrations, and reaction driving forces which is probabilistically most in agreement with experimentally measured metabolite concentrations (Methods). Using this framework, we computed a maximum-likelihood thermodynamic state for the model, simulating aerobic growth on glucose. In the thermodynamic state computed by PMO, we found that all metabolite concentrations lie within physiologically reasonable ranges (1 $\mu$M – 1 mM). In addition, all “anomalous concentrations” identified by the framework (metabolite concentrations lying more than one standard deviation away from the mean experimental value) had been identified and explained previously [gollub_probabilistic_2021], and can potentially be addressed by lumping together those reactions for which substrate channelling is known to happen, as reported in literature.
@@
We used the PMO-derived thermodynamic state to identify candidate bottlenecks, in terms of enzyme demand, across the network. To this end, we first computed the flux-force efficacy of each reaction in the thermodynamic state [noor_pathway_2014] (Methods). The flux-force efficacy is a unitless quantity, ranging from 00 to $1$, denoting the ratio between net flux (forward minus backward flux) and total flux (forward plus backward flux) of a reaction. Reactions operating at low flux-force efficacy have a lower net flux due to two reasons: (1) the forward flux is lower in absolute terms, and (2) the higher backward flux is counter-productive and subtracts from the forward flux. Therefore, to achieve a given required net flux, the cell has to invest more resources in maintaining a higher enzyme level. To identify potential thermodynamic bottlenecks that lead to high enzyme costs in specific reactions, we therefore screened all reactions for their predicted flux-force efficacy and flux (Figure 7B) and identified those predicted to operate at low efficacy, while carrying significant flux (Figure 7B, labelled points).
To assess to which extent the thermodynamic operating states predicted by PMO are predictive of enzyme investment, we split reactions into two groups based on whether their predicted flux-force efficacy sits above or below $50\%$, and compared the distributions of measured enzyme abundances between the two groups (Methods, Supplementary Figure S20). While there are many determinants of enzyme abundance beyond thermodynamics, including enzyme turnover, affinity, or regulation, we observed a significant difference between the two distributions ($p<0.001$, two-sided Wilcoxon rank-sum test), with the low efficacy having approximately a 3-fold higher median enzyme abundance than the high efficacy group. Establishing a causal link can be very difficult, but we can speculate that this negative correlation might be explained by metabolism evolving to compensate for the low efficacy of some reactions with a higher expression level of their catalysing enzyme [noor_pathway_2014, khana_thermodynamics_2025].
@@
Figure 7: Thermodynamic analysis of the model via the curated thermodynamic parameter set. A Probabilistic max-min driving force (MDF) analysis of flux distributions obtained by parsimonious flux balance analysis for a total of $12$ growth conditions. All flux distributions tested have a positive MDF, implying they are thermodynamically feasible under reasonable physiological metabolite concentration ranges. The computed MDF values cluster in three groups, corresponding to glycolytic aerobic, gluconeogenic aerobic, and anaerobic growth conditions. B: Fluxes relative to the glucose uptake flux (EX_glc__D_e) and flux-force efficacies computed by probabilistic metabolic optimisation (PMO). The labelled data points represent examples of reactions (excluding transport and spontaneous reactions) with low predicted flux-force efficacy (here, below $20\%$), but carrying high relative flux in the optimal solution (here, more than $5\%$ of the glucose uptake flux). xyl__D: D-xylose; glc__D: D-glucose; rib__D: D-ribose; glyc: glycerol; akg: alpha-keto-glutarate; ac: acetate; pyr: pyruvate; lac__D: D-lactate; succ: succinate. fum: fumarate.